# Repeated Padding for Sequential Recommendation

### v2

@@ 3.5. Implementation Details
Table 4. Performance improvement on the five short-sequence datasets. The ‘w/ RP’ represents adding our method. All improvements are statistically significant, as determined by a paired t-test with $p\leq 0.05$.
Dataset Metric GRU4Rec (Hidasi et al., 2015b) w/ RP Improve NARM (Li et al., 2017) w/ RP Improve Caser (Tang and Wang, 2018) w/ RP Improve NextItNet (Yuan et al., 2019) w/ RP Improve
Toys Hit@5 0.0134 0.0276 105.97% 0.0275 0.0286 4.00% 0.0132 0.0152 15.15% 0.0229 0.0273 19.21%
Hit@10 0.0256 0.0434 69.53% 0.0401 0.0415 3.49% 0.0231 0.0246 6.49% 0.0391 0.0452 15.60%
Hit@20 0.0432 0.0652 50.93% 0.0598 0.0589 -1.51% 0.0364 0.0380 4.40% 0.0539 0.0613 13.73%
NDCG@5 0.0083 0.0179 115.66% 0.0179 0.0188 5.03% 0.0079 0.0096 21.52% 0.0128 0.0153 19.53%
NDCG@10 0.0121 0.0229 89.26% 0.0220 0.0233 5.91% 0.0111 0.0126 13.51% 0.0172 0.0201 16.86%
NDCG@20 0.0166 0.0284 71.08% 0.0269 0.0276 2.60% 0.0149 0.0160 7.38% 0.0217 0.0251 15.67%
Beauty Hit@5 0.0182 0.0286 57.14% 0.0213 0.0244 14.55% 0.0149 0.0199 33.56% 0.0223 0.0263 17.94%
Hit@10 0.0327 0.0468 43.12% 0.0367 0.0400 8.99% 0.0261 0.0330 26.44% 0.0393 0.0442 12.47%
Hit@20 0.0551 0.0732 32.85% 0.0617 0.0659 6.81% 0.0457 0.0511 11.82% 0.0633 0.0725 14.53%
NDCG@5 0.0116 0.0175 50.86% 0.0131 0.0152 16.03% 0.0090 0.0127 41.11% 0.0135 0.0151 11.85%
NDCG@10 0.0162 0.0234 44.44% 0.0180 0.0202 12.22% 0.0126 0.0169 34.13% 0.0191 0.0221 15.71%
NDCG@20 0.0219 0.0300 36.99% 0.0243 0.0258 6.17% 0.0175 0.0214 22.29% 0.0251 0.0280 11.55%
Sports Hit@5 0.0078 0.0114 46.15% 0.0126 0.0140 11.11% 0.0081 0.0092 13.58% 0.0104 0.0112 7.69%
Hit@10 0.0143 0.0204 42.66% 0.0218 0.0238 9.17% 0.0146 0.0160 9.59% 0.0195 0.0215 10.26%
Hit@20 0.0249 0.0338 35.74% 0.0361 0.0385 6.65% 0.0257 0.0265 3.11% 0.0338 0.0368 8.88%
NDCG@5 0.0044 0.0073 65.91% 0.0076 0.0088 15.79% 0.0048 0.0056 16.67% 0.0064 0.0074 15.63%
NDCG@10 0.0065 0.0100 53.85% 0.0106 0.0120 13.21% 0.0069 0.0078 13.04% 0.0093 0.0103 10.75%
NDCG@20 0.0091 0.0134 47.25% 0.0142 0.0157 10.56% 0.0096 0.0105 9.38% 0.0129 0.0139 7.75%
Yelp Hit@5 0.0092 0.0119 29.35% 0.0140 0.0159 13.57% 0.0131 0.0144 9.92% 0.0109 0.0121 11.01%
Hit@10 0.0172 0.0223 29.65% 0.0255 0.0277 8.63% 0.0238 0.0252 5.88% 0.0218 0.0240 10.09%
Hit@20 0.0314 0.0407 29.62% 0.0443 0.0482 8.80% 0.0414 0.0435 5.07% 0.0400 0.0433 8.25%
NDCG@5 0.0057 0.0072 26.32% 0.0084 0.0098 16.67% 0.0084 0.0089 5.95% 0.0075 0.0081 8.00%
NDCG@10 0.0083 0.0105 26.51% 0.0120 0.0136 13.33% 0.0118 0.0127 7.63% 0.0106 0.0117 10.38%
NDCG@20 0.0118 0.0151 27.97% 0.0167 0.0188 12.57% 0.0162 0.0173 6.79% 0.0149 0.0162 8.72%
Home Hit@5 0.0030 0.0063 110.00% 0.0059 0.0073 23.73% 0.0035 0.0044 25.71% 0.0039 0.0043 10.26%
Hit@10 0.0057 0.0108 89.47% 0.0109 0.0125 14.68% 0.0066 0.0075 13.64% 0.0074 0.0085 14.86%
Hit@20 0.0106 0.0182 71.70% 0.0185 0.0205 10.81% 0.0125 0.0136 8.80% 0.0137 0.0151 10.22%
NDCG@5 0.0018 0.0040 122.22% 0.0037 0.0048 29.73% 0.0022 0.0026 18.18% 0.0025 0.0029 16.00%
NDCG@10 0.0027 0.0054 100.00% 0.0053 0.0065 22.64% 0.0031 0.0039 25.81% 0.0037 0.0042 13.51%
NDCG@20 0.0039 0.0073 87.18% 0.0072 0.0085 18.06% 0.0046 0.0050 8.70% 0.0051 0.0059 15.69%
Dataset Metric SASRec (Kang and McAuley, 2018) w/ RP Improve LightSANs (Fan et al., 2021) w/ RP Improve FMLP-Rec (Zhou et al., 2022) w/ RP Improve STOSA (Fan et al., 2022) w/ RP Improve
Toys Hit@5 0.0472 0.0587 24.36% 0.0376 0.0388 3.19% 0.0418 0.0463 10.77% 0.0486 0.0611 25.72%
Hit@10 0.0716 0.0843 17.74% 0.0505 0.0564 11.68% 0.0570 0.0632 10.88% 0.0699 0.0865 23.75%
Hit@20 0.0965 0.1142 18.34% 0.0656 0.0770 17.38% 0.0759 0.0852 12.25% 0.0976 0.1196 22.54%
NDCG@5 0.0317 0.0393 23.97% 0.0265 0.0260 -1.89% 0.0289 0.0314 8.65% 0.0330 0.0416 26.06%
NDCG@10 0.0396 0.0476 20.20% 0.0307 0.0316 2.93% 0.0338 0.0368 8.88% 0.0398 0.0498 25.13%
NDCG@20 0.0471 0.0551 16.99% 0.0345 0.0368 6.67% 0.0386 0.0423 9.59% 0.0467 0.0581 24.41%
Beauty Hit@5 0.0409 0.0477 16.63% 0.0312 0.0328 5.13% 0.0387 0.0439 13.44% 0.0387 0.0468 20.93%
Hit@10 0.0637 0.0726 13.97% 0.0469 0.0485 3.41% 0.0566 0.0648 14.49% 0.0613 0.0708 15.50%
Hit@20 0.0942 0.1034 9.77% 0.0709 0.0741 4.51% 0.0814 0.0905 11.18% 0.0909 0.1047 15.18%
NDCG@5 0.0257 0.0310 20.62% 0.0201 0.0217 7.96% 0.0260 0.0286 10.00% 0.0248 0.0313 26.21%
NDCG@10 0.0331 0.0390 17.82% 0.0255 0.0267 4.71% 0.0317 0.0353 11.36% 0.0320 0.0390 21.88%
NDCG@20 0.0408 0.0468 14.71% 0.0319 0.0323 1.25% 0.0380 0.0418 10.00% 0.0394 0.0476 20.81%
Sports Hit@5 0.0187 0.0237 26.74% 0.0130 0.0150 15.38% 0.0149 0.0201 34.90% 0.0218 0.0271 24.31%
Hit@10 0.0299 0.0369 23.41% 0.0208 0.0231 11.06% 0.0238 0.0294 23.53% 0.0325 0.0429 32.00%
Hit@20 0.0447 0.0553 23.71% 0.0327 0.0352 7.65% 0.0345 0.0439 27.25% 0.0482 0.0640 32.78%
NDCG@5 0.0124 0.0154 24.19% 0.0087 0.0097 11.49% 0.0097 0.0132 36.08% 0.0146 0.0174 19.18%
NDCG@10 0.0160 0.0196 22.50% 0.0112 0.0122 8.93% 0.0126 0.0162 28.57% 0.0181 0.0225 24.31%
NDCG@20 0.0197 0.0242 22.84% 0.0142 0.0153 7.75% 0.0153 0.0199 30.07% 0.0220 0.0277 25.91%
Yelp Hit@5 0.0168 0.0210 25.00% 0.0120 0.0159 32.50% 0.0121 0.0154 27.27% 0.0164 0.0216 31.71%
Hit@10 0.0282 0.0359 27.30% 0.0207 0.0267 28.99% 0.0209 0.0258 23.44% 0.0283 0.0367 29.68%
Hit@20 0.0443 0.0587 32.51% 0.0346 0.0429 23.99% 0.0356 0.0425 19.38% 0.0471 0.0593 25.90%
NDCG@5 0.0106 0.0129 21.70% 0.0073 0.0099 35.62% 0.0070 0.0099 41.43% 0.0103 0.0134 30.10%
NDCG@10 0.0142 0.0177 24.65% 0.0101 0.0134 32.67% 0.0098 0.0133 35.71% 0.0141 0.0183 29.79%
NDCG@20 0.0183 0.0234 27.87% 0.0136 0.0175 28.68% 0.0135 0.0175 29.63% 0.0189 0.0240 26.98%
Home Hit@5 0.0096 0.0130 35.42% 0.0059 0.0065 10.17% 0.0079 0.0085 7.59% 0.0115 0.0138 20.00%
Hit@10 0.0155 0.0204 31.61% 0.0100 0.0098 -2.00% 0.0125 0.0128 2.40% 0.0171 0.0196 14.62%
Hit@20 0.0244 0.0311 27.46% 0.0173 0.0185 6.94% 0.0193 0.0191 -1.04% 0.0262 0.0312 19.08%
NDCG@5 0.0061 0.0087 42.62% 0.0034 0.0039 14.71% 0.0051 0.0055 7.84% 0.0072 0.0077 6.94%
NDCG@10 0.0083 0.0116 39.76% 0.0047 0.0050 6.38% 0.0066 0.0069 4.55% 0.0092 0.0101 9.78%
NDCG@20 0.0101 0.0137 35.64% 0.0064 0.0067 4.69% 0.0083 0.0085 2.41% 0.0115 0.0130 13.04%
Table 5. Performance improvement on the long-sequence dataset. The ‘w/ RP’ represents adding our method.
Dataset Metric GRU4Rec (Hidasi et al., 2015b) w/ RP Improve NARM (Li et al., 2017) w/ RP Improve Caser (Tang and Wang, 2018) w/ RP Improve NextItNet (Yuan et al., 2019) w/ RP Improve
ML-1M Hit@5 0.1104 0.1407 27.45% 0.0493 0.0541 9.74% 0.0465 0.0495 6.45% 0.0922 0.1008 9.33%
Hit@10 0.1841 0.2200 19.50% 0.0868 0.0879 1.27% 0.0823 0.0841 2.19% 0.1491 0.1623 8.85%
Hit@20 0.2849 0.3300 15.83% 0.1459 0.1470 0.75% 0.1416 0.1429 0.92% 0.2354 0.2519 7.01%
NDCG@5 0.0704 0.0914 29.83% 0.0297 0.0349 17.51% 0.0277 0.0304 9.75% 0.0620 0.0643 3.71%
NDCG@10 0.0941 0.1168 24.12% 0.0417 0.0457 9.59% 0.0398 0.0414 4.02% 0.0815 0.0866 6.26%
NDCG@20 0.1195 0.1446 21.00% 0.0565 0.0606 7.26% 0.0541 0.0562 3.88% 0.0963 0.1025 6.44%
Metric SASRec (Kang and McAuley, 2018) w/ RP Improve LightSANs (Fan et al., 2021) w/ RP Improve FMLP-Rec (Zhou et al., 2022) w/ RP Improve STOSA (Fan et al., 2022) w/ RP Improve
[…]

### v3

@@ 4.1.4. Implementation Details
Table 4. Performance improvement on five short-sequence datasets. The ‘w/ RP’ and ‘w/ RP+’ represent adding RepPad and Our RepPad+ to the corresponding model, respectively. The best performance of the three is bolded. The ‘Imp-O’ and ‘Imp-R’ represent the relative improvements of RepPad+ over the original model and the RepPad, respectively. All improvements are statistically significant, as determined by a paired t-test with $p\leq 0.05$.
Dataset Metric GRU4Rec w/ RP w/ RP+ Imp-O Imp-R NARM w/ RP w/ RP+ Imp-O Imp-R Caser w/ RP w/ RP+ Imp-O Imp-R NextItNet w/ RP w/ RP+ Imp-O Imp-R
Toys H@5 0.0134 0.0276 0.0319 138.06% 15.58% 0.0275 0.0286 0.0309 12.36% 8.04% 0.0132 0.0152 0.0173 31.06% 13.82% 0.0229 0.0273 0.0292 27.51% 6.96%
H@10 0.0256 0.0434 0.0482 88.28% 11.06% 0.0401 0.0415 0.0439 9.48% 5.78% 0.0231 0.0246 0.0269 16.45% 9.35% 0.0391 0.0452 0.0463 18.41% 2.43%
H@20 0.0432 0.0652 0.0716 65.74% 9.82% 0.0598 0.0589 0.0611 2.17% 3.74% 0.0364 0.0380 0.0403 10.71% 6.05% 0.0539 0.0613 0.0630 16.88% 2.77%
N@5 0.0083 0.0179 0.0217 161.45% 21.23% 0.0179 0.0188 0.0208 16.20% 10.64% 0.0079 0.0096 0.0109 37.97% 13.54% 0.0128 0.0153 0.0161 25.78% 5.23%
N@10 0.0121 0.0229 0.0269 122.31% 17.47% 0.0220 0.0233 0.0251 14.09% 7.73% 0.0111 0.0126 0.0145 30.63% 15.08% 0.0172 0.0201 0.0219 27.33% 8.96%
N@20 0.0166 0.0284 0.0328 97.59% 15.49% 0.0269 0.0276 0.0294 9.29% 6.52% 0.0149 0.0160 0.0177 18.79% 10.63% 0.0217 0.0251 0.0263 21.20% 4.78%
Beauty H@5 0.0182 0.0286 0.0306 68.13% 6.99% 0.0213 0.0244 0.0283 32.86% 15.98% 0.0149 0.0199 0.0215 44.30% 8.04% 0.0223 0.0263 0.0269 20.63% 2.28%
H@10 0.0327 0.0468 0.0505 54.43% 7.91% 0.0367 0.0400 0.0447 21.80% 11.75% 0.0261 0.0330 0.0340 30.27% 3.03% 0.0393 0.0442 0.0456 16.03% 3.17%
H@20 0.0551 0.0732 0.0756 37.21% 3.28% 0.0617 0.0659 0.0682 10.53% 3.49% 0.0457 0.0511 0.0525 14.88% 2.74% 0.0633 0.0725 0.0743 17.38% 2.48%
N@5 0.0116 0.0175 0.0198 70.69% 13.14% 0.0131 0.0152 0.0187 42.75% 23.03% 0.0090 0.0127 0.0139 54.44% 9.45% 0.0135 0.0151 0.0157 16.30% 3.97%
N@10 0.0162 0.0234 0.0263 62.35% 12.39% 0.0180 0.0202 0.0240 33.33% 18.81% 0.0126 0.0169 0.0179 42.06% 5.92% 0.0191 0.0221 0.0235 23.04% 6.33%
N@20 0.0219 0.0300 0.0326 48.86% 8.67% 0.0243 0.0258 0.0294 20.99% 13.95% 0.0175 0.0214 0.0223 27.43% 4.21% 0.0251 0.0280 0.0289 15.14% 3.21%
Sports H@5 0.0078 0.0114 0.0133 70.51% 16.67% 0.0126 0.0140 0.0157 24.60% 12.14% 0.0081 0.0092 0.0101 24.69% 9.78% 0.0104 0.0112 0.0123 18.27% 9.82%
H@10 0.0143 0.0204 0.0232 62.24% 13.73% 0.0218 0.0238 0.0249 14.22% 4.62% 0.0146 0.0160 0.0165 13.01% 3.13% 0.0195 0.0215 0.0234 20.00% 8.84%
H@20 0.0249 0.0338 0.0383 53.82% 13.31% 0.0361 0.0385 0.0398 10.25% 3.38% 0.0257 0.0265 0.0272 5.84% 2.64% 0.0338 0.0368 0.0386 14.20% 4.89%
N@5 0.0044 0.0073 0.0083 88.64% 13.70% 0.0076 0.0088 0.0103 35.53% 17.05% 0.0048 0.0056 0.0068 41.67% 21.43% 0.0064 0.0074 0.0079 23.44% 6.76%
N@10 0.0065 0.0100 0.0115 76.92% 15.00% 0.0106 0.0120 0.0132 24.53% 10.00% 0.0069 0.0078 0.0089 28.99% 14.10% 0.0093 0.0103 0.0110 18.28% 6.80%
N@20 0.0091 0.0134 0.0153 68.13% 14.18% 0.0142 0.0157 0.0170 19.72% 8.28% 0.0096 0.0105 0.0116 20.83% 10.48% 0.0129 0.0139 0.0146 13.18% 5.04%
Yelp H@5 0.0092 0.0119 0.0149 61.96% 25.21% 0.0140 0.0159 0.0164 17.14% 3.14% 0.0131 0.0144 0.0148 12.98% 2.78% 0.0109 0.0121 0.0128 17.43% 5.79%
H@10 0.0172 0.0223 0.0274 59.30% 22.87% 0.0255 0.0277 0.0296 16.08% 6.86% 0.0238 0.0252 0.0269 13.03% 6.75% 0.0218 0.0240 0.0252 15.60% 5.00%
H@20 0.0314 0.0407 0.0474 50.96% 16.46% 0.0443 0.0482 0.0505 14.00% 4.77% 0.0414 0.0435 0.0503 21.50% 15.63% 0.0400 0.0433 0.0451 12.75% 4.16%
N@5 0.0057 0.0072 0.0089 56.14% 23.61% 0.0084 0.0098 0.0100 19.05% 2.04% 0.0084 0.0089 0.0087 3.57% -2.25% 0.0075 0.0081 0.0086 14.67% 6.17%
N@10 0.0083 0.0105 0.0129 55.42% 22.86% 0.0120 0.0136 0.0142 18.33% 4.41% 0.0118 0.0127 0.0134 13.56% 5.51% 0.0106 0.0117 0.0125 17.92% 6.84%
N@20 0.0118 0.0151 0.0179 51.69% 18.54% 0.0167 0.0188 0.0194 16.17% 3.19% 0.0162 0.0173 0.0184 13.58% 6.36% 0.0149 0.0162 0.0166 11.41% 2.47%
Home H@5 0.0030 0.0063 0.0073 143.33% 15.87% 0.0059 0.0073 0.0078 32.20% 6.85% 0.0035 0.0044 0.0047 34.29% 6.82% 0.0039 0.0043 0.0046 17.95% 6.98%
H@10 0.0057 0.0108 0.0121 112.28% 12.04% 0.0109 0.0125 0.0135 23.85% 8.00% 0.0066 0.0075 0.0080 21.21% 6.67% 0.0074 0.0085 0.0089 20.27% 4.71%
H@20 0.0106 0.0182 0.0201 89.62% 10.44% 0.0185 0.0205 0.0216 16.76% 5.37% 0.0125 0.0136 0.0146 16.80% 7.35% 0.0137 0.0151 0.0160 16.79% 5.96%
N@5 0.0018 0.0040 0.0047 161.11% 17.50% 0.0037 0.0048 0.0052 40.54% 8.33% 0.0022 0.0026 0.0029 31.82% 11.54% 0.0025 0.0029 0.0032 28.00% 10.34%
N@10 0.0027 0.0054 0.0063 133.33% 16.67% 0.0053 0.0065 0.0070 32.08% 7.69% 0.0031 0.0039 0.0040 29.03% 2.56% 0.0037 0.0042 0.0045 21.62% 7.14%
N@20 0.0039 0.0073 0.0083 112.82% 13.70% 0.0072 0.0085 0.0090 25.00% 5.88% 0.0046 0.0050 0.0056 21.74% 12.00% 0.0051 0.0059 0.0063 23.53% 6.78%
Average - - - 84.11% 14.85% - - - 20.86% 8.38% - - - 24.24% 8.17% - - - 19.03% 5.57%
Dataset Metric SASRec w/ RP w/ RP+ Imp-O Imp-R LightSANs w/ RP w/ RP+ Imp-O Imp-R gMLP w/ RP w/ RP+ Imp-O Imp-R FMLP-Rec w/ RP w/ RP+ Imp-O Imp-R
Toys H@5 0.0472 0.0587 0.0622 31.78% 5.96% 0.0376 0.0388 0.0438 16.49% 12.89% 0.0373 0.0385 0.0394 5.63% 2.34% 0.0418 0.0463 0.0473 13.16% 2.16%
H@10 0.0716 0.0843 0.0866 20.95% 2.73% 0.0505 0.0564 0.0614 21.58% 8.87% 0.0495 0.0508 0.0523 5.66% 2.95% 0.0570 0.0632 0.0660 15.79% 4.43%
H@20 0.0965 0.1142 0.1172 21.45% 2.63% 0.0656 0.0770 0.0843 28.51% 9.48% 0.0665 0.0685 0.0712 7.07% 3.94% 0.0759 0.0852 0.0899 18.45% 5.52%
N@5 0.0317 0.0393 0.0433 36.59% 10.18% 0.0265 0.0260 0.0308 16.23% 18.46% 0.0269 0.0276 0.0285 5.95% 3.26% 0.0289 0.0314 0.0333 15.22% 6.05%
N@10 0.0396 0.0476 0.0512 29.29% 7.56% 0.0307 0.0316 0.0364 18.57% 15.19% 0.0308 0.0321 0.0333 8.12% 3.74% 0.0338 0.0368 0.0393 16.27% 6.79%
N@20 0.0471 0.0551 0.0589 25.05% 6.90% 0.0345 0.0368 0.0422 22.32% 14.67% 0.0351 0.0359 0.0378 7.69% 5.29% 0.0386 0.0423 0.0454 17.62% 7.33%
Beauty H@5 0.0409 0.0477 0.0512 25.18% 7.34% 0.0312 0.0328 0.0398 27.56% 21.34% 0.0313 0.0361 0.0355 13.42% -1.66% 0.0387 0.0439 0.0459 18.60% 4.56%
H@10 0.0637 0.0726 0.0762 19.62% 4.96% 0.0469 0.0485 0.0598 27.51% 23.30% 0.0486 0.0525 0.0539 10.91% 2.67% 0.0566 0.0648 0.0664 17.31% 2.47%
H@20 0.0942 0.1034 0.1089 15.61% 5.32% 0.0709 0.0741 0.0854 20.45% 15.25% 0.0703 0.0748 0.0765 8.82% 2.27% 0.0814 0.0905 0.0935 14.86% 3.31%
N@5 0.0257 0.0310 0.0344 33.85% 10.97% 0.0201 0.0217 0.0269 33.83% 23.96% 0.0209 0.0239 0.0244 16.75% 2.09% 0.0260 0.0286 0.0305 17.31% 6.64%
N@10 0.0331 0.0390 0.0424 28.10% 8.72% 0.0255 0.0267 0.0334 30.98% 25.09% 0.0264 0.0292 0.0314 18.94% 7.53% 0.0317 0.0353 0.0371 17.03% 5.10%
N@20 0.0408 0.0468 0.0507 24.26% 8.33% 0.0319 0.0323 0.0398 24.76% 23.22% 0.0319 0.0348 0.0361 13.17% 3.74% 0.0380 0.0418 0.0439 15.53% 5.02%
Sports H@5 0.0187 0.0237 0.0294 57.22% 24.05% 0.0130 0.0150 0.0173 33.08% 15.33% 0.0136 0.0152 0.0169 24.26% 11.18% 0.0149 0.0201 0.0209 40.27% 3.98%
H@10 0.0299 0.0369 0.0434 45.15% 17.62% 0.0208 0.0231 0.0280 34.62% 21.21% 0.0210 0.0229 0.0248 18.10% 8.30% 0.0238 0.0294 0.0307 28.99% 4.42%
H@20 0.0447 0.0553 0.0641 43.40% 15.91% 0.0327 0.0352 0.0417 27.52% 18.47% 0.0328 0.0342 0.0365 11.28% 6.73% 0.0345 0.0439 0.0481 39.42% 9.57%
N@5 0.0124 0.0154 0.0194 56.45% 25.97% 0.0087 0.0097 0.0114 31.03% 17.53% 0.0090 0.0101 0.0112 24.44% 10.89% 0.0097 0.0132 0.0129 32.99% -2.27%
N@10 0.0160 0.0196 0.0239 49.38% 21.94% 0.0112 0.0122 0.0148 32.14% 21.31% 0.0114 0.0126 0.0137 20.18% 8.73% 0.0126 0.0162 0.0157 24.60% -3.09%
N@20 0.0197 0.0242 0.0291 47.72% 20.25% 0.0142 0.0153 0.0182 28.17% 18.95% 0.0143 0.0154 0.0167 16.78% 8.44% 0.0153 0.0199 0.0206 34.64% 3.52%
Yelp H@5 0.0168 0.0210 0.0219 30.36% 4.29% 0.0120 0.0159 0.0163 35.83% 2.52% 0.0108 0.0115 0.0140 29.63% 21.74% 0.0121 0.0154 0.0213 76.03% 38.31%
H@10 0.0282 0.0359 0.0376 33.33% 4.74% 0.0207 0.0267 0.0280 35.27% 4.87% 0.0190 0.0201 0.0239 25.79% 18.91% 0.0209 0.0258 0.0319 52.63% 23.64%
H@20 0.0443 0.0587 0.0598 34.99% 1.87% 0.0346 0.0429 0.0437 26.30% 1.86% 0.0325 0.0338 0.0416 28.00% 23.08% 0.0356 0.0425 0.0467 31.18% 9.88%
N@5 0.0106 0.0129 0.0136 28.30% 5.43% 0.0073 0.0099 0.0103 41.10% 4.04% 0.0066 0.0072 0.0085 28.79% 18.06% 0.0070 0.0099 0.0144 105.71% 45.45%
N@10 0.0142 0.0177 0.0188 32.39% 6.21% 0.0101 0.0134 0.0139 37.62% 3.73% 0.0093 0.0099 0.0116 24.73% 17.17% 0.0098 0.0133 0.0178 81.63% 33.83%
N@20 0.0183 0.0234 0.0247 34.97% 5.56% 0.0136 0.0175 0.0180 32.35% 2.86% 0.0126 0.0134 0.0161 27.78% 20.15% 0.0135 0.0175 0.0215 59.26% 22.86%
Home H@5 0.0096 0.0130 0.0136 41.67% 4.62% 0.0059 0.0065 0.0085 44.07% 30.77% 0.0057 0.0080 0.0083 45.61% 3.75% 0.0079 0.0085 0.0110 39.24% 29.41%
H@10 0.0155 0.0204 0.0211 36.13% 3.43% 0.0100 0.0098 0.0133 33.00% 35.71% 0.0092 0.0123 0.0126 36.96% 2.44% 0.0125 0.0128 0.0165 32.00% 28.91%
H@20 0.0244 0.0311 0.0324 32.79% 4.18% 0.0173 0.0185 0.0202 16.76% 9.19% 0.0157 0.0181 0.0193 22.93% 6.63% 0.0193 0.0191 0.0243 25.91% 27.23%
N@5 0.0061 0.0087 0.0090 47.54% 3.45% 0.0034 0.0039 0.0057 67.65% 46.15% 0.0037 0.0054 0.0057 54.05% 5.56% 0.0051 0.0055 0.0092 80.39% 67.27%
N@10 0.0083 0.0116 0.0124 49.40% 6.90% 0.0047 0.0050 0.0072 53.19% 44.00% 0.0049 0.0067 0.0071 44.90% 5.97% 0.0066 0.0069 0.0090 36.36% 30.43%
N@20 0.0101 0.0137 0.0145 43.56% 5.84% 0.0064 0.0067 0.0089 39.06% 32.84% 0.0065 0.0082 0.0086 32.31% 4.88% 0.0083 0.0085 0.0109 31.33% 28.24%
Average - - - 35.34% 8.92% - - - 31.25% 18.10% - - - 21.29% 8.03% - - - 34.99% 15.37%
Table 5. Performance improvement on two long-sequence datasets. The ‘w/ RP’ and ‘w/ RP+’ represent adding RepPad and RepPad+ to the corresponding model, respectively. The best performance of the three is bolded. The ‘Imp-O’ and ‘Imp-R’ represent the relative improvements of RepPad+ over the original model and the RepPad, respectively.
Dataset Metric GRU4Rec w/ RP w/ RP+ Imp-O Imp-R NARM w/ RP w/ RP+ Imp-O Imp-R Caser Imp-OA Imp-RP Imp-O Imp-R NextItNet w/ RP w/ RP+ Imp-O Imp-R
LastFM H@5 0.0229 0.0330 0.0353 54.15% 6.97% 0.0505 0.0578 0.0596 18.02% 3.11% 0.0321 0.0394 0.0426 32.71% 8.12% 0.0312 0.0339 0.0367 17.63% 8.26%
H@10 0.0339 0.0541 0.0582 71.68% 7.58% 0.0716 0.0862 0.0890 24.30% 3.25% 0.0550 0.0679 0.0707 28.55% 4.12% 0.0560 0.0565 0.0679 21.25% 20.18%
H@20 0.0624 0.0706 0.0869 39.26% 23.09% 0.1037 0.1220 0.1312 26.52% 7.54% 0.0862 0.0908 0.1055 22.39% 16.19% 0.0855 0.0927 0.1138 33.10% 22.76%
N@5 0.0152 0.0225 0.0243 59.87% 8.00% 0.0341 0.0374 0.0400 17.30% 6.95% 0.0226 0.0292 0.0308 36.28% 5.48% 0.0192 0.0215 0.0243 26.56% 13.02%
N@10 0.0188 0.0292 0.0307 63.30% 5.14% 0.0408 0.0465 0.0494 21.08% 6.24% 0.0300 0.0383 0.0396 32.00% 3.39% 0.0272 0.0285 0.0341 25.37% 19.65%
N@20 0.0259 0.0333 0.0385 48.65% 15.62% 0.0490 0.0556 0.0599 22.24% 7.73% 0.0380 0.0442 0.0470 23.68% 6.33% 0.0348 0.0376 0.0457 31.32% 21.54%
ML-1M H@5 0.1104 0.1407 0.1492 35.14% 6.04% 0.0493 0.0541 0.0532 7.91% -1.66% 0.0465 0.0495 0.0507 9.03% 2.42% 0.0922 0.1008 0.1052 14.10% 4.37%
H@10 0.1841 0.2200 0.2313 25.64% 5.14% 0.0868 0.0879 0.0923 6.34% 5.01% 0.0823 0.0841 0.0862 4.74% 2.50% 0.1491 0.1623 0.1674 12.27% 3.14%
H@20 0.2849 0.3300 0.3482 22.22% 5.52% 0.1459 0.1470 0.1548 6.10% 5.31% 0.1416 0.1429 0.1501 6.00% 5.04% 0.2354 0.2519 0.2600 10.45% 3.22%
N@5 0.0704 0.0914 0.0942 33.81% 3.06% 0.0297 0.0349 0.0360 21.21% 3.15% 0.0277 0.0304 0.0310 11.91% 1.97% 0.0620 0.0643 0.0657 5.97% 2.18%
N@10 0.0941 0.1168 0.1207 28.27% 3.34% 0.0417 0.0457 0.0464 11.27% 1.53% 0.0398 0.0414 0.0426 7.04% 2.90% 0.0815 0.0866 0.0883 8.34% 1.96%
N@20 0.1195 0.1446 0.1502 25.69% 3.87% 0.0565 0.0606 0.0628 11.15% 3.63% 0.0541 0.0562 0.0572 5.73% 1.78% 0.0963 0.1025 0.1063 10.38% 3.71%
Average - - - 42.31% 7.78% - - - 16.12% 4.32% - - - 18.34% 5.02% - - - 18.06% 10.33%
Dataset Metric SASRec w/ RP w/ RP+ Imp-O Imp-R LightSANs w/ RP w/ RP+ Imp-O Imp-R gMLP w/ RP w/ RP+ Imp-O Imp-R FMLP-Rec w/ RP w/ RP+ Imp-O Imp-R
LastFM H@5 0.0440 0.0505 0.0569 29.32% 12.67% 0.0422 0.0303 0.0385 -8.77% 27.06% 0.0239 0.0247 0.0349 46.03% 41.30% 0.0257 0.0321 0.0376 46.30% 17.13%
H@10 0.0606 0.0789 0.0826 36.30% 4.69% 0.0560 0.0560 0.0679 21.25% 21.25% 0.0385 0.0422 0.0541 40.52% 28.20% 0.0431 0.0459 0.0532 23.43% 15.90%
H@20 0.0972 0.1220 0.1284 32.10% 5.25% 0.0789 0.0817 0.0936 18.63% 14.57% 0.0670 0.0695 0.0862 28.66% 24.03% 0.0560 0.0661 0.0752 34.29% 13.77%
N@5 0.0287 0.0336 0.0373 29.97% 11.01% 0.0274 0.0186 0.0233 -14.96% 25.27% 0.0140 0.0150 0.0241 72.14% 60.67% 0.0170 0.0208 0.0214 25.88% 2.88%
N@10 0.0340 0.0430 0.0454 33.53% 5.58% 0.0318 0.0267 0.0327 2.83% 22.47% 0.0186 0.0217 0.0304 63.44% 40.09% 0.0224 0.0253 0.0259 15.63% 2.37%
N@20 0.0434 0.0538 0.0567 30.65% 5.39% 0.0377 0.0333 0.0393 4.24% 18.02% 0.0257 0.0275 0.0384 49.42% 39.64% 0.0256 0.0304 0.0315 23.05% 3.62%
ML-1M H@5 0.1748 0.1674 0.1818 4.00% 8.60% 0.1179 0.1096 0.1295 9.84% 18.16% 0.0601 0.0742 0.0868 44.43% 16.98% 0.1045 0.1109 0.1328 27.08% 19.75%
H@10 0.2624 0.2608 0.2715 3.47% 4.10% 0.1873 0.1758 0.2026 8.17% 15.24% 0.1070 0.1180 0.1323 23.64% 12.12% 0.1667 0.1639 0.1851 11.04% 12.93%
H@20 0.3762 0.3780 0.3839 2.05% 1.56% 0.2765 0.2647 0.3003 8.61% 13.45% 0.1717 0.1882 0.1998 16.37% 6.16% 0.2454 0.2445 0.2609 6.32% 6.71%
N@5 0.1171 0.1096 0.1173 0.17% 7.03% 0.0770 0.0712 0.0842 9.35% 18.26% 0.0368 0.0483 0.0551 49.73% 14.08% 0.0670 0.0719 0.0829 23.73% 15.30%
N@10 0.1450 0.1396 0.1462 0.83% 4.73% 0.0993 0.0924 0.1078 8.56% 16.67% 0.0518 0.0625 0.0697 34.56% 11.52% 0.0871 0.0889 0.1025 17.68% 15.30%
N@20 0.1737 0.1693 0.1745 0.46% 3.07% 0.1218 0.1153 0.1323 8.62% 14.74% 0.0680 0.0800 0.0867 27.50% 8.38% 0.1069 0.1092 0.1216 13.75% 11.36%
Average - - - 16.90% 6.14% - - - 6.36% 18.76% - - - 41.37% 25.26% - - - 22.35% 11.42%
4.2. Overall Improvements
[…]